OpenStudy (anonymous):

How to simplify (9/4)i?

OpenStudy (jdoe0001):

well get the factors for 9 then the factors for 4 say \(\bf \cfrac{9}{4}\implies \cfrac{3\cdot 3\cdot 1}{2\cdot 2\cdot 1}\) so.. .none of the factors atop cancel out with the bottom factors, so, \(\cfrac{9}{4}\) simplified is \(\bf \cfrac{9}{4}\)